### v5

# Dynamic tariffs-based demand response in retail electricity market under uncertainty

Abstract
Demand response (DR) programs play a crucial role in improving system reliability and mitigating price volatility by altering the core profile of electricity consumption. This paper proposes a game-theoretical model that captures the dynamic interplay between retailers (leaders) and consumers (followers) in a tariffs-based electricity market under uncertainty. The proposed procedure offers theoretical and economic insights by analyzing demand flexibility within a hierarchical decision-making framework. In particular, two main market configurations are examined under uncertainty: i) there exists a retailer that exercises market power over consumers, and ii) the retailer and the consumers participate in a perfect competitive game. The former case is formulated as a mathematical program with equilibrium constraints (MPEC), whereas the latter case is recast as a mixed-integer linear program (MILP). These problems are solved by deriving equivalent tractable reformulations based on the Karush-Kuhn-Tucker (KKT) optimality conditions of each agent’s problem.
Numerical simulations based on real data from the European Energy Exchange platform are used to illustrate the performance of the proposed methodology. The results indicate that the proposed model effectively characterizes the interactions between retailers and flexible consumers in both perfect and imperfect market structures. Under perfect competition, the economic benefits extend not only to consumers but also to overall social welfare. Conversely, in an imperfect market, retailers leverage consumer flexibility to enhance their expected profits, transferring the risk of uncertainty to end-users. Additionally, the degree of consumer flexibility and their valuation of electricity consumption play significant roles in shaping market outcomes. These findings highlight the crucial impact of market structure and consumer behavior on the dynamics of electricity market pricing under demand response programs.
keywords:
Bi-level programming, Tariff-based demand response, Market power, MILP, MPEC, Retail-consumer equilibrium.
journal: a Journal
1 Introduction
[…]
One potential solution is demand response (DR), which is typically achieved through incentive-based or price-based consumer flexibility. These methods offer consumers financial incentives to reduce or shift their peak loads to off-peak periods. To that end, consumers facing a high price during a given hour can adopt one of the following strategies (or a combination of them): i) reducing consumption, ii) shifting consumption to low-price periods, and/or iii) covering part of the consumption by on-site distributive generation or storage sources. This behavior can be used to soften peak loads at given hours and reduce electricity market prices caused by marginal carbon-based technologies (Nolan and O’Malley, 2015; Guo and Weeks, 2022). Moreover, since DR is a distributed resource located at the end of the distribution system, additional environmental benefits are obtained from a reduction of electricity losses in the transmission and distribution lines (Shen et al., 2014). Thus, the main advantages of DR programs are: i) they can contribute to reduce system costs and manage price volatility in real-time (Zugno et al., 2013; Aussel et al., 2020; Alipour et al., 2019), ii) they allow higher penetration of intermittent RES in the electric power system (Morales et al., 2013; Hakimi et al., 2019; Antunes et al., 2020), iii) they help improve system security by contributing to the real-time balance between generation and the demand (Zugno et al., 2013; Wang et al., 2020), and iv) they can mitigate electricity retailers’ market power by increasing consumers flexibility (Devine and Bertsch, 2023).
DR programs employ bidirectional communication to send price signals and financial incentives to responsive consumers, as well as demand signals to retailers or demand aggregators. Upon receiving updated information regarding price tariffs, consumers can calculate their benefits and the trade-off of shifting their consumption at critical times for system operators. The literature on DR is extensive and covers different incentive-based strategies, such as offering consumers dynamic electricity tariffs like time-of-use (TOU), critical peak pricing (CPP), and real-time pricing (RTP). Dynamic tariffs can be attractive to retailers as they allow them to align retail price tariffs with spot market prices, thus transferring some of the retailer-centered risks to customers (Guo and Weeks, 2022; Askeland et al., 2020; Clastres and Khalfallah, 2021). Using tariffs-based DR, Askeland et al. (2020) formulate the electricity grid tariff design problem with a bilevel structure in the context of prosumers at the end-user level. They address the tariff optimization problem as a tool for indirect load control and analyze how a simple tariff scheme can be used to activate end-user flexibility and efficiently reduce grid load by developing Mathematical Problems with Equilibrium Constraints (MPECs). Applying a bi-level programming problem, Clastres and Khalfallah (2021) study the impact of dynamic pricing and load-shifting on a retailer’s electricity supply to consumer markets. Considering the consumer benefit function, Guo and Weeks (2022) model a dynamic price-based DR and find that implementing demand-side management stimuli can reduce consumer electricity demand and further increase retail profit.
However, significant gaps persist in the literature as the integration of RES increases and the demand for consumer flexibility: i) incorporating the characterization of uncertainties affecting electricity prices, production costs, and demands in the day-ahead (DA) market. ii) explicitly studying consumer behavior and leveraging their flexibility for operational and planning decisions. iii) comparing both perfect and imperfect electricity markets, involving multiple participants with varying market powers, to analyze both theoretical and practical economic implications.
Overall, the increase in energy costs, the soaring global competition for fuel, and the tightening of emission requirements place huge pressure on power systems. This motivates us to extend the potential of DR programs to enhance system reliability, ensure supply security, mitigate market power, and manage price volatility. These benefits are achieved while accounting for uncertainty across various market configurations.
The objective of this work is to define meaningful decision-making tools and to analyze potential market interactions between electricity retailers and responsive consumers participating in the day-ahead electricity market. Specifically, we seek to mimic the hierarchy of the decision-making process, where a retailer needs to set first the electricity price tariffs for the consumers, and then they would react accordingly. Moreover, motivated by the concept of smart grids, which promotes different forms of horizontal consumer aggregations (cooperatives, energy communities, virtual power plants, etc.), we assume two market configurations. In the first one, the retailer has market power by anticipating the reaction of consumers, which is modeled by a bilevel (Stackelberg) problem. In the second one, we assume that the retailer is myopic to the consumers’ response, which is modeled with a competitive equilibrium. Furthermore, for both cases and extending previous works, we improve the characterization of the consumers’ demand response by using a (non-linear) quadratic utility function.
These facts, together with accounting for different levels of market competition and sources of uncertainty, have a great impact on the retailers’ and consumption decision-making processes, and ultimately on the resulting market outcomes. Therefore, this work models the profit maximization problems of retailers and the utility maximization of consumers in electricity consumption, taking into account the inherent uncertainties present in both perfect and imperfect electricity markets. Specifically, the retailer’s problem with market power is formulated as a bi-level program and solved as an MPEC (the interested reader is referred to Luo et al. (1996) for a complete treatment of bilevel programs and their solution methods), and the equilibrium model with perfect competition is solved as a mixed-integer linear programming (MILP) (the interested reader is referred to Leyffer and Munson (2010) for a complete treatment of how equilibrium problems can be solved as mixed complementarity problems (MCP)). Uncertainty is dealt with stochastic programming techniques, which extends the majority of deterministic DR models available in the literature.
The paper presents several key findings: i) it demonstrates that retailers can maximize their expected profits by exercising market power and utilizing consumer flexibility to mitigate risk. ii) it shows that consumers can minimize their procurement costs, and increase overall social welfare under a perfect equilibrium model. iii) the proposed models are versatile and capable of adapting to various groups of consumers with flexible demand across different market configurations. iv) explicitly modeling consumer behavior through a utility function in a dynamic pricing environment provides deeper insights into characterizing end-users. This approach reveals how their flexibility can aid the integration of renewable energy sources into the power system under DR programs. A key takeaway is the increased vulnerability of the electricity market to the market power of oligopolistic entities when demand-side flexibility is neglected, leading to market inefficiencies. However, harnessing consumer flexibility in operational and strategic planning decisions can counterbalance retailers’ market power and mitigate grid instability.
The remainder of this paper is organized as follows. Section 2 reviews the current literature on the topic. Section 3 demonstrates the model formulation which entails the retailer’s and consumers’ problems. The formulation of the problems as an MPEC, and linear reformulation of the problem as MILP in the equilibrium are presented. Computational and simulation results are reported and discussed in Section 4. Finally, Section 5 closes by drawing some concluding remarks from the provided discussions and results.
2 Literature review
From the supply side, power producers, encompassing both conventional and RES, sell their generated power to the grid through organized energy markets. These markets are responsible for handling transactions related to energy and ancillary services. On the demand side, DR programs play an integral role. These programs, managed by DR buyers such as load-serving entities, distribution system operators (DSOs), large consumers, and DR aggregators, are pivotal in implementing strategies to collect and manage flexibility options from end-users. This management effectively modifies energy consumption patterns in response to supply conditions or market signals.
DR buyers, while engaging in various market activities, primarily focus on interacting with consumers through contracts or DR programs to create system flexibility. Such flexibility is particularly crucial in day-ahead and intraday markets, where supply and demand adjustments occur closer to real-time operations. The specific nature and extent of these interactions and transactions can vary, depending on the particular market design and regulatory environment.
In the literature, DR programs have been considered pivotal in addressing a range of economic and operational challenges. These challenges include achieving economic efficiency, reducing market volatility, optimizing utility in electricity consumption, minimizing discomfort in existing consumption patterns, enhancing grid stability, optimizing system performance, and reducing the carbon footprint. Studies such as those by Devine and Bertsch (2023); Guo and Weeks (2022); Askeland et al. (2020); Clastres and Khalfallah (2021); Nilsson et al. (2018), among others have extensively explored these facets. From a retail perspective, the importance of dynamic tariffs-based DR is particularly noteworthy. These tariffs align retail prices with spot market prices, thereby transferring some of the risks centered around retailers to the customers. As highlighted by Nilsson et al. (2018), dynamic tariffs are recognized as one of the most effective strategies to enhance demand flexibility. This strategy not only complements the DR programs but also serves as a critical link between fluctuating market prices and consumer behavior.
Furthermore, the role of DR in general, and specifically in mitigating market power, is examined by Devine and Bertsch (2023) using a market equilibrium model. This study underscores the importance of distinguishing between regular consumers and prosumers with distributed energy resources (DERs). Such a distinction is crucial for a quantitative analysis of the effects and benefits of DR programs, illustrating the nuanced ways in which DR programs can influence market dynamics and consumer participation.
The demand response literature, from a theoretical perspective, has made significant strides in addressing market complexities. It encompasses studies considering imperfect markets using bilevel models, such as Afşar et al. (2016); Soares et al. (2019); Aussel et al. (2020); Soares et al. (2020) and perfect competition with equilibrium models such as Devine and Bertsch (2023, 2018); Hu et al. (2016); Nguyen et al. (2016).
Afşar et al. (2016) present a bi-level optimization model where the supplier establishes time-differentiated electricity prices, and the consumers minimize their electricity consumption costs to enhance grid efficiency. Soares et al. (2019) present a model considering a single leader (retailer) whose aim is to maximize profit and multi-followers (consumers) whose objective is to minimize their electricity cost. Aussel et al. (2020) propose the interactions among electricity suppliers, local agents, aggregators, and consumers with a trilevel multi-leader-multi-follower game for load shifting induced by the time of use pricing. Soares et al. (2020) develop a comprehensive model in which a retailer maximizes its profit and a cluster of consumers react to the decision of the retailer based on electricity prices by determining the operation of the controllable loads to minimize the electricity bill and a monetized discomfort factor associated with the indoor temperature deviations.
Moreover, DR programs have been extensively analyzed using equilibrium models in various electricity market applications. The widespread application of these models in electricity market research stems from their robust ability to represent diverse market structures and the interactions between market participants. For instance, Devine and Bertsch (2023) investigate the role of DR in mitigating market power, particularly through load shifting and self-generation. Devine and Bertsch (2018) delves into the costs associated with providing flexibility through load shedding and self-generation using DERs. Nguyen et al. (2016) analyzes the energy scheduling challenges faced by load-serving entities, particularly in managing flexible and inflexible loads. Additionally, Hu et al. (2016) present a stochastic-multi-objective Nash-Cournot model, which aims to reduce peak demand, energy prices, and emissions.
Furthermore, the application of tariff-based DR strategies has been explored in various contexts. For distribution network (Lu et al., 2018), optimal bidding strategy in day-ahead electricity markets (Vayá and Andersson, 2014), optimal pricing strategy in pool-based electricity markets (Ruiz and Conejo, 2009), modeling coordinated cyber-physical attacks (Li et al., 2015), and optimal charging schedules of plug-in electric vehicles (Momber et al., 2015), among others. Summarizing the benefits of DR programs from the perspective of various electricity market participants is a complex task, given the extensive range of literature on the subject. However, these benefits have been summarized in Figure 1.
Consumers System Reliability - Incentive payments - System optimization - Improved comfort - Bill savings - Grid stability - Reduced volatility - Customer participation - Diversified resources Electricity Market Environment - Deferred infrastructure costs - Price reduction - Capacity increase - Reduced market power - High renewable penetration - Rational use of energy resources - High economic efficiency - Carbon footprint reduction
Figure 1: Enhanced summary of DR programs benefits.
The existing literature on dynamic electricity price tariffs extensively examines their impact on individual players in the energy market. This includes areas such as cost minimization for demand aggregators, optimal reconfiguration of microgrids by DSOs, optimal allocation of distributed generations by TSOs, retailer profit maximization, and consumer benefit enhancement. Notable studies in this domain include those by Ajoulabadi et al. (2020); Wang et al. (2020); Nejad et al. (2019); Dagoumas and Polemis (2017); Nilsson et al. (2018), among others. The task of explicitly modeling consumer behavior in response to dynamic pricing is particularly challenging in this field. This entails understanding their price elasticities or utility functions, particularly during uncertain times. Such modeling remains a demanding and highly complex research topic in DR programs. Addressing different market configurations, including both strategic and myopic retailers, and explicitly modeling consumer behavior through utility functions further complicates this task. The integration of stochastic and non-dispatchable RES makes it difficult to manage these systems efficiently.
By leveraging the advantages of game-theoretical models to address market power and competitive markets, in this work, we develop a tariffs-based demand response program by simultaneously solving retailers’ and consumers’ problems. Assuming that consumers are rational decision-makers and risk-averse to their electricity consumption, we explicitly model their behavior, considering utility functions to quantify their welfare. While measuring satisfaction is an abstract concept, in classical microeconomics, individuals’ behavior is analyzed by the choices they make when purchasing a product, which is represented by utility functions or indifference curves (Rubinstein et al., 2006).
Thus, the main contributions of this paper are:
  1. i)
    Application of demand response to address retailer and consumer decision-making challenges in electricity markets with dynamic pricing. Imperfect competition is analyzed using a Stackelberg game and perfect competition is through equilibrium models, incorporating uncertainties in RES and demand. The focus is on optimizing social welfare and mitigating market power through the strategic use of consumer flexibility.
  2. ii)
    The retailer-consumer market power interaction is transformed into a single-level problem using a tractable MPEC. On the other hand, the equilibrium is addressed as an MILP problem, where consumer demand is treated as a variable for both upper- and lower-level problems, and dynamic pricing is considered an exogenous parameter. In addition, an NLP reformulation is employed, providing solutions that are equivalent to those of the MILP for the equilibrium problem.
  3. iii)
    Introducing a utility function to quantify consumer benefits from flexibility, which, together with the modeling of battery storage, extends consumers’ flexibility to two possibilities: i) reacting to the price tariff where they can decide to buy more or less energy at a particular hour, and ii) delaying or anticipating consumption within hours. Modeling these two flexibilities has significant operational and economic implications.
  4. iv)
    Utilizing an extensive set of real-world case studies to illustrate the effectiveness of our model in capturing market structures and the implications of DR programs on the operational and economic decisions of market participants. The numerical results focus on analyzing the impact of consumer flexibility on welfare and other market outcomes. The insights derived from this analysis can be useful, not only for market participants and regulators but also to guide future research in this field.
3 Problem formulation
Most electricity markets exhibit a hierarchical decision-making structure, that allows participants with market power to potentially anticipate the reactions of subsequent decision-making entities. In this regard, bilevel programming is an adequate optimization-based modeling technique to represent this sort of interaction. It can be defined in various forms, including single-leader single-follower, single-leader multi-follower, multi-leader single-follower, and multi-leader multi-follower games. While the single leader-single follower case has been extensively analyzed since von Stackelberg’s seminal work (Von Stackelberg, 1952), its extensions, accounting for multiple followers, remain less explored. Additionally, there aren’t general numerical methods for finding equilibrium solutions efficiently, and the existing algorithms are problem-specific.
Table 1: Main notations used in the model formulation for quick reference.
Notations Description
Sets
$J$ set of consumers ranging from $j=1,...,\mid J\mid$
$T$ hours considered ranging from $t=1,...,|T|$
$\Omega$ set of scenarios ranging from $\omega=1,...,\mid\Omega\mid$
Parameters
$A_{jt\omega}$ linear coefficient of utility function for consumer $j$, hour $t$, scenario $\omega$ (€/kWh)
[…]
$C$ penalization cost for power imbalances (€/kWh)
$\Delta_{j}^{C_{max}}$ level of flexibility of consumer $j$ (kWh)
Variables
$\Delta_{jt\omega}^{C}$ variation of consumer’s demand $j$, hour $t$, scenario $\omega$ (kWh)
$q_{jt\omega}$ electricity purchased by consumer $j$, hour $t$, scenario $\omega$ (kWh)
$q_{t\omega}^{S}$ electricity purchased by retailer in the spot market, hour $t$, scenario $\omega$ (kWh)
$\delta_{t\omega}$ electricity imbalance by retailer at hour $t$, scenario $\omega$ [kWh]
$P_{t}$ retailer price tariff at hour $t$ (€/kWh)
$y_{t\omega}$ electricity imbalance by retailer at hour $t$, scenario $\omega$ [kWh]
$P_{t\omega}^{S}$ spot price at hour $t$, scenario $\omega$ (€/kWh)
Dual Variables
$\alpha_{t\omega}^{+}/\alpha_{t\omega}^{-}$ upper/lower limits for electricity imbalance for retailer hour $t$, scenario $\omega$
$\beta_{t\omega}$ nonnegative electricity imbalance at hour $t$, scenario $\omega$
$\lambda_{j\omega}$ total quantity variation of demand for consumer $j$ at scenario $\omega$
$\varepsilon_{jt\omega}$ nonnegativity flexible and inflexible consumption of consumer $j$ at hour $t$, at scenario $\omega$
$\mu_{t\omega}$ electricity imbalance hour $t$, scenario $\omega$
$\nu_{jt\omega}^{C_{max}}/\nu_{jt\omega}^{C_{min}}$ maximum/minimum flexibility for consumer $j$ at hour $t$, scenario $\omega$
$\theta_{t\omega}$ nonnegative spot market electricity at hour $t$, scenario $\omega$
$\gamma_{t}$ nonnegative tariff decided by retailer at hour $t$
In this work, we consider a single retailer who wants to optimize its expected profit, and $\left|J\right|$ consumers who want to minimize the difference between the cost of purchasing electricity and the utility brought by their consumption. That is, the retailer determines the selling prices to maximize its expected profit, which depends on the consumers’ purchases and the consumers determine their load demands, which depend on the retail price tariffs. The consumers want to minimize the disutility costs. First, we develop a bi-level programming problem assuming the retailers behave as leaders (upper-level) and the consumers as followers (lower-level). Then, for comparison purposes, we analyze a competitive equilibrium between the retailer and consumers, assuming no market power.
The bi-level programming with a strategic retailer, is reformulated as an MPEC problem. This is done by recasting the consumers’ problems using their KKT optimality conditions, which are introduced as constraints in the retailer’s problem. Uncertainties related to consumers’ demand, spot market quantities, and prices are addressed using stochastic programming.
Even though electricity generation costs are approximated by quadratic production costs, consumer utilities are frequently simplified with linear functions. Quadratic utility functions are interesting in DR programs because of their non-decreasing marginal benefits (Niromandfam et al., 2020; Chen et al., 2010). To demonstrate that the overall implications of the proposed models are not solely dependent on a specific form of the utility function, we employ both the linear utility function and its quadratic extension, which is also a popular form for describing customer behavior using elasticity (Niromandfam et al., 2020).
Due to the assumption that the model provides no possibility for consumers to switch to a different retailer (short-term model), the utility function prevents the retailer from increasing the price tariffs possibly up to infinity to maximize its profit [^1] To handle such challenges, for example, Zugno et al. (2013) include minimum, maximum, and average prices into the consumers’ problem as constraints..
For the equilibrium model with perfect competition, we formulate a MILP problem by linearizing the complementarity conditions of the corresponding retailer and consumers’ KKT optimality conditions. In addition, an NLP reformulation of the equilibrium problem is derived using the methods proposed by Leyffer and Munson (2010), which provides a solution equivalent to the MILP formulation.
The model has first-stage and second-stage decisions. Prior to the day-ahead market clearance, the retailer decides the price tariff for hour $t$ (first stage). However, the spot market price, the quantity purchased by the retailer from the electricity market, the quantity bought by consumers, and the electricity quantity imbalance decisions are second-stage and scenario-dependent variables.
3.1 Retailer’s problem
The retailer maximizes her expected profit subject to technical and economic constraints, and accounting for uncertainty, which is modeled by a discrete number of scenarios. The retailer’s problem is stated mathematically as follows:
$\displaystyle\max_{\bf{X}}$ $\sum_{\omega\in{\Omega}}\sigma_{\omega}\left(\sum_{j\in{J},t\in{T}}P_{t}q_{jt\omega}-\sum_{t\in{T}}(P_{t\omega}^{S}q_{t\omega}^{S}+Cy_{t\omega})\right)$ (3.1a)
subject to
$\sum_{j\in{J}}q_{jt\omega}-q_{t\omega}^{S}=\delta_{t\omega}\quad:\mu_{t\omega}\quad\forall t,\forall\omega$ (3.1b)
$\delta_{t\omega}\leq y_{t\omega}\quad:\alpha_{t\omega}^{+}\quad\forall t,\forall\omega$ (3.1c)
$-\delta_{t\omega}\leq y_{t\omega}\quad:\alpha_{t\omega}^{-}\quad\forall t,\forall\omega$ (3.1d)
$y_{t\omega}\geq 0\quad:\beta_{t\omega}\quad\forall t,\forall\omega$ (3.1e)
$q_{t\omega}^{S}\geq 0\quad:\theta_{t\omega}\quad\forall t,\forall\omega$ (3.1f)
$P_{t}\geq 0\quad:\gamma_{t}\quad\forall t$ (3.1g)
where $\bf{X}$=$\{P_{t},q_{t\omega}^{S},\delta_{t\omega},y_{t\omega}\}$ is the retailer’s set of decision variables, while $\mu_{t\omega},\alpha_{t\omega}^{+},\alpha_{t\omega}^{-}$, $\theta_{t\omega},\gamma_{t},\beta_{t\omega}$ are the dual variables associated with the constraints ((3.1b))-((3.1g)).
The first term in the objective function is the revenue for the retailer from the quantity bought by consumer $j$, at hour $t$, with scenario $\omega$ and price $P_{t}$, which is the retail price of electricity (tariff), a first-stage decision determined by the retailer for each hour $t$. The second term is the cost of purchasing the electricity quantity $q_{t\omega}^{S}$ at the spot market price $P_{t\omega}^{S}$ at hour $t$ and with scenario $\omega$. The last term is a cost (penalization of electricity imbalances), which can be positive or negative (variable $y_{t\omega}$ represents the absolute value of deviations of the electricity imbalances), and $\sigma_{\omega}$ is the probability associated with scenario $\omega$. The summation of the product over index $\omega\in\Omega$ represents the expected profit, which is what the retailer wants to maximize over the time horizon.
Constraint ((3.1b)) is the electricity balancing constraint that guarantees the total demand by all consumers at hour $t$ with scenario $\omega$ minus retailer’s supply is the imbalance at that particular time. The linear set of constraints ((3.1c)), ((3.1d)) and ((3.1e)) ensure that $y_{t\omega}=|\delta_{t\omega}|$. These imbalances are penalized in the objective function with weight $C$. Constraints ((3.1f)) and ((3.1g)) are non-negativity constraints on the spot market quantity and price decided by the retailer for tariff at hour $t$, respectively. $P_{t}$ is a first-stage decision, which needs to be settled before the uncertainty realization, while $q_{jt\omega},q_{t\omega}^{S},\delta_{t\omega}$, $P_{t\omega}^{S}$ and $y_{t\omega}$ are second stage decisions. Correspondingly, the dual variables for each group of constraints are indicated on the right side of a colon.
3.2 Consumers’ problem
In modeling the consumers’ problem with DR, it is common to express the type of consumers (residential, industrial, or commercial) explicitly so that part of their consumption is considered flexible, and part of the consumption is inelastic (demand that must be met strictly at all times). For instance, Halvgaard et al. (2012) model heat pumps for heating residential buildings, where the heating system of the house becomes a flexible consumption in the smart grid. Zugno et al. (2013) extend the above model by considering the heating dynamics of a building, indoor temperature, or temperature inside a water tank where flexibility (the output of interest) is characterized through the indoor temperature. They also consider the inflexible part of consumption in their model. However, in our model setting, including individual appliances would require a large number of new variables and constraints, which will further complicate its numerical resolution and the economical interpretation of the results that we aspire to deliver. Moreover, we believe that the main demand patterns/profiles can be already accounted for by the appropriate inter-temporal adjustment of the utility function parameters.
By considering the consumers’ utility function explicitly in the objective function, we explore a residential consumers’ model with flexibility towards the dynamic price reported by the retailer. Thus, the consumers face an economic problem where it is crucial to balance the trade-off between the cost of electricity procurement and the discomfort of deviating from the utility of consumption given a flexible price and the possibility to shift the consumption to low price periods. In particular, the consumers’ problem can be represented within a game-theoretical setting as it is parametrized by the decision of the retailer (hourly price $P_{t}$). The problem for consumer $j$ with scenario $\omega$ can be expressed as:
$\displaystyle\min_{q_{jt\omega},\Delta_{jt\omega}^{C}}$ $\sum_{t\in{T}}\left(P_{t}q_{jt\omega}-A_{jt\omega}(q_{jt\omega}+\Delta_{jt\omega}^{C})+\frac{1}{2}B_{jt\omega}(q_{jt\omega}+\Delta_{jt\omega}^{C})^{2}\right)$ (3.2a)
subject to
$q_{jt\omega}+\Delta_{jt\omega}^{C}\geq 0\quad:\varepsilon_{jt\omega}\quad\forall j,\forall t,\forall\omega$ (3.2b)
$-\Delta_{j}^{C_{max}}\leq\Delta_{jt\omega}^{C}\leq+\Delta_{j}^{C_{max}}\quad:\nu_{jt\omega}^{C_{min}},\nu_{jt\omega}^{C_{max}}\quad\forall j,\forall t,\forall\omega$ (3.2c)
$\sum_{t\in{T}}\Delta_{jt\omega}^{C}=0\quad:\lambda_{j\omega}\quad\forall j,\forall\omega$ (3.2d)
where the objective function is the minus social welfare (cost minus utility) and $q_{jt\omega}$ (inelastic demand) and $\Delta_{jt\omega}^{C}$ (flexible demand) are consumer’s decision variables. In the objective function $P_{t}q_{jt\omega}$ is the cost of purchasing $q_{jt\omega}$ electricity and $A_{jt\omega}(q_{jt\omega}+\Delta_{jt\omega}^{C})-\frac{1}{2}B_{jt\omega}(q_{jt\omega}+\Delta_{jt\omega}^{C})^{2}$ is the quadratic utility function that measures the benefit consumer $j$ achieves by consuming the amount of energy ($q_{jt\omega}+\Delta_{jt\omega}^{C}$) during hour $t$, scenario $\omega$.
Considering the consumers’ objective function without any further constraints, consumption would only take place in periods when the real-time price is lower than the marginal benefit. Note that the marginal utility is a demand function $P_{t}=A_{jt\omega}-B_{jt\omega}(q_{jt\omega}+\Delta_{jt\omega}^{C})$ where electricity quantity demanded decreases as electricity price increases. In other words, at the optimum the marginal utility (benefit) from consuming electricity equals the retail price tariff $P_{t}$.
Parameters $A_{jt\omega}$ and $B_{jt\omega}$ are the intercept and slope of the demand function, which are important parameters to characterize the behavior of consumers toward the market outcomes of the model. Different values for these parameters can capture the dynamics of consumers’ demand. Constraint ((3.2c)) sets the lower and upper limits for flexible consumption, where $\Delta_{j}^{C_{max}}$ is the level of flexibility of consumer $j$. The inequalities guarantee that the flexible part of consumption at hour $t$ falls in the range between $-\Delta_{j}^{C_{max}}$ and $+\Delta_{j}^{C_{max}}$, which bound consumption increase and decrease, respectively. If $\Delta_{j}^{C_{max}}=0$, there is no flexibility in demand, while $\Delta_{j}^{C_{max}}>0$, corresponds to the flexible demand. In the latter case, the consumer considers a dynamic price tariff that ensures cost savings and better welfare. Besides, constraint ((3.2d)) ensures the net total flexible demand during the planning time horizon is zero. In the consumers’ problem, $q_{jt\omega}$ and $\Delta_{jt\omega}^{C}$ are second-stage decisions. Note that since the dynamic electricity price $P_{t}$ enters the consumers’ problem as a parameter (it is only a variable in the retailer problem), the optimization problems of the consumers are convex (quadratic objective function with linear constraints). Therefore, we can replace each consumer’s problem with its corresponding KKT optimality conditions (Haghighat and Kennedy, 2012), which are sufficient for optimality.
3.2.1 KKT formulation of the consumer problem
The KKT optimality conditions of problem ((3.2)) are:
$\frac{\partial\mathcal{L}_{j\omega}}{\partial q_{jt\omega}}=P_{t}-A_{jt\omega}+B_{jt\omega}(q_{jt\omega}+\Delta_{jt\omega}^{C})-\varepsilon_{jt\omega}=0\quad\forall j,\forall t,\forall\omega$ (3.3a)
$\frac{\partial\mathcal{L}_{j\omega}}{\partial\Delta_{jt\omega}^{C}}=-A_{jt\omega}+B_{jt\omega}(q_{jt\omega}+\Delta_{jt\omega}^{C})-\nu_{jt\omega}^{C_{min}}+\nu_{jt\omega}^{C_{max}}-\lambda_{j\omega}-\varepsilon_{jt\omega}=0\quad\forall j,\forall t,\forall\omega$ (3.3b)
$\sum_{t\in{T}}\Delta_{jt\omega}^{C}=0\quad\forall j,\forall\omega$ (3.3c)
$0\leq q_{jt\omega}+\Delta_{jt\omega}^{C}\perp\varepsilon_{jt\omega}\geq 0\quad\forall j,\forall t,\forall\omega$ (3.3d)
$0\leq\Delta_{jt\omega}^{C}+\Delta_{j}^{C_{max}}\perp\nu_{jt\omega}^{C_{min}}\geq 0\quad\forall j,\forall t,\forall\omega$ (3.3e)
$0\leq+\Delta_{j}^{C_{max}}-\Delta_{jt\omega}^{C}\perp\nu_{jt\omega}^{C_{max}}\geq 0\quad\forall j,\forall t,\forall\omega$ (3.3f)
$\nu_{jt\omega}^{C_{min}},\nu_{jt\omega}^{C_{max}},\varepsilon_{jt\omega}\geq 0\quad\forall j,\forall t,\forall\omega$ (3.3g)
$\lambda_{j\omega}\quad\text{free}\quad\forall j,\forall\omega$ (3.3h)
where $\mathcal{L}_{j\omega}$ stands for the Lagrangian function for consumer $j$ and scenario $\omega$ of problem ((3.2)). By convention, the symbol $\perp$ indicates complementarity so that any of the two inequalities is satisfied as equality, i.e., the product of each expression and the corresponding dual variable must be zero. Note that the system of KKT optimality conditions is linear, with the exception of the complementarity conditions ((3.3d))-((3.3f)).
3.3 The Retailer’s MPEC problem
In game theory, hierarchical optimization problems of this type can be formulated mathematically as MPECs. Figure 2 depicts the modeling framework for the MPEC problem where the leader optimizes his economic problem subject to his constraints, and by considering the equilibrium among the followers.
Figure 2: Game-theoretical framework for modeling demand response as a single leader-followers problem. Solid arrows indicate information signals from the retailer, while dashed arrows represent inferences on the consumers’ behavior.
Thus, the MPEC is expressed as follows:
$\displaystyle\min_{\bf{X}^{MPEC}}$ $-\sum_{\omega\in{\Omega}}\sigma_{\omega}\left(\sum_{t\in{T},j\in{J}}P_{t}q_{jt\omega}-\sum_{t\in{T}}(P_{t\omega}^{S}q_{t\omega}^{S}+Cy_{t\omega})\right)$ (3.4a)
subject to
[…]
where $\bf{X}^{MPEC}$=$\{P_{t},q_{t\omega}^{S},\delta_{t\omega},y_{t\omega},q_{jt\omega},\Delta_{jt\omega}^{c},\nu_{jt\omega}^{C_{max}},\nu_{jt\omega}^{C_{min}},\varepsilon_{jt\omega},\lambda_{j\omega}\}$ is the set of retailer’s decision variables.
The MPEC is thus formulated as a single-level optimization problem, where the retailer minimizes his expected negative profit subject to other constraints and the KKT optimality conditions of the lower-level problems. However, the MPEC model in ((3.3)) has two sources of nonlinearities: the bilinear product $P_{t}q_{jt\omega}$ in the objective function, and the complementarity conditions ((3.3d))-((3.3f)) in the consumers’ KKT optimality conditions.
The second type of nonlinearities can be linearized using the strategy proposed in Fortuny-Amat and McCarl (1981), which is by applying the big-M technique. However, we still face the nonlinear product $P_{t}q_{jt\omega}$ which, as opposed to other problem settings Ruiz and Conejo (2009) and Sadat and Fan (2017) cannot be linearized by employing the strong duality theorem. However, due to the observed good numerical performance, we have opted to solve the original MPEC problem with a state-of-the-art NLP solver like KNITRO (Artelys KNITRO, 2021).
3.4 The equilibrium model
Considering the perfect competition, we tackle the equilibrium model by jointly solving retailers’ and consumers’ KKT optimality conditions using two techniques: i) by linearizing and casting the KKT systems of equations as a MILP problem employing the Fortuny-Amat and McCarl (1981) big-M technique, and ii) by directly addressing and reformulating the problem as a NLP using the approach in Leyffer and Munson (2010). However, we mainly present the reformulation of the equilibrium problem as a MILP and its results, for our analysis. For details on the reformulation of NLP problems, the interested reader is referred to Abate et al. (2021).
3.4.1 KKT formulation of the retailer problem
To derive the retailer’s KKT optimality conditions and to deal with the equilibrium model, we assume that the market is in perfect competition, where players compete in terms of deciding their optimal quantities. Thus, $q_{jt\omega}$ also becomes a decision variable for the retailer. This implies that price tariff $P_{t}$ is considered exogenous for both retailers’ and consumers’ problems, and its equilibrium value will naturally come out as the market clearing price, i.e., the price that satisfies all the players’ (retailers’ and consumers’) optimality conditions.
The KKT optimality conditions derived from the retailer’s problem are:
$\frac{\partial\mathcal{L}_{\omega}}{\partial q_{jt\omega}}=\sigma_{\omega}P_{t}+\mu_{t\omega}=0\quad\forall j,\forall t,\forall\omega$ (3.5a)
$\frac{\partial\mathcal{L}_{\omega}}{\partial q_{t\omega}^{S}}=\mu_{t\omega}+\sigma_{\omega}P_{t\omega}^{S}-\theta_{t\omega}=0\quad\forall t,\forall\omega$ (3.5b)
$\frac{\partial\mathcal{L}_{\omega}}{\partial\delta_{t\omega}}=\mu_{t\omega}+\alpha_{t\omega}^{+}-\alpha_{t\omega}^{-}=0\quad\forall t,\forall\omega$ (3.5c)
$\frac{\partial\mathcal{L}_{\omega}}{\partial y_{t\omega}}=\sigma_{\omega}C-\alpha_{t\omega}^{+}-\alpha_{t\omega}^{-}-\beta_{t\omega}=0\quad\forall t,\forall\omega$ (3.5d)
$\sum_{j\in{J}}q_{jt\omega}-q_{t\omega}^{S}=\delta_{t\omega}\quad\forall t,\forall\omega$ (3.5e)
$0\leq y_{t\omega}-\delta_{t\omega}\perp\alpha_{t\omega}^{+}\geq 0\quad\forall t,\forall\omega$ (3.5f)
$0\leq y_{t\omega}+\delta_{t\omega}\perp\alpha_{t\omega}^{-}\geq 0\quad\forall t,\forall\omega$ (3.5g)
$0\leq q_{t\omega}^{S}\perp\theta_{t\omega}\geq 0\quad\forall t,\forall\omega$ (3.5h)
$0\leq y_{t\omega}\perp\beta_{t\omega}\geq 0\quad\forall t,\forall\omega$ (3.5i)
$\alpha_{t\omega}^{+},\alpha_{t\omega}^{-},\beta_{t\omega},\theta_{t\omega}\geq 0\quad\forall t,\forall\omega$ (3.5j)
$\mu_{t\omega}\quad\text{free}\quad\forall t,\forall\omega.$ (3.5k)
3.5 MILP formulation
We solve the equilibrium problem as MILP by concatenating the KKT optimality conditions (((3.3a))-((3.3h)) and ((3.5a))-((3.5k))). The complementarity slackness conditions ((3.3d))-((3.3f)) and ((3.5f))-((3.5i)) are linearized by introducing an auxiliary variable for each condition.
Then to exploit the efficient performance of MILP solvers, the resulting system of MIL constraints can be reformulated as a MILP by including an arbitrary constant as the objective function:
$\displaystyle\min_{\bf{X}^{Comp}}1$ (3.6a)
subject to
[…]
$\mu_{t\omega}\quad\text{free}\quad\forall t,\forall\omega$ (3.6g)
where $\bf{X}^{Comp}$=$\{q_{t\omega}^{S},\delta_{t\omega},y_{t\omega},q_{jt\omega},\Delta_{jt\omega}^{C},\nu_{jt\omega}^{C_{min}},\nu_{jt\omega}^{C_{max}},\varepsilon_{jt\omega},\lambda_{j\omega},\mu_{t\omega},\alpha_{t\omega}^{+},\alpha_{t\omega}^{-},\theta_{t\omega},\beta_{t\omega}$,
$\psi_{jt\omega},\psi_{jt\omega}^{C_{min}},\psi_{jt\omega}^{C_{max}},\tau_{t\omega}^{+},\tau_{t\omega}^{-},\tau_{t\omega}^{q},\tau_{t\omega}^{y},\tau_{t}^{P}\}$ is the set of the decision variables for the competitive model. In particular, the complementarity conditions from the consumer’s KKT conditions can be linearized as follows:
$\Delta_{jt\omega}^{C}+\Delta_{j}^{C_{max}}\geq 0,\quad\forall j,\forall t,\forall\omega$ (3.7a)
$\Delta_{j}^{C_{max}}-\Delta_{jt\omega}^{C}\geq 0,\quad\forall j,\forall t,\forall\omega$ (3.7b)
$q_{jt\omega}+\Delta_{jt\omega}^{C}\geq 0\quad\forall j,\forall t,\forall\omega$ (3.7c)
$\nu_{jt\omega}^{C_{min}}\geq 0\quad\forall j,\forall t,\forall\omega$ (3.7d)
$\nu_{jt\omega}^{C_{max}}\geq 0\quad\forall j,\forall t,\forall\omega$ (3.7e)
$\varepsilon_{jt\omega}\geq 0\quad\forall j,\forall t,\forall\omega$ (3.7f)
$\Delta_{jt\omega}^{C}+\Delta_{j}^{C_{max}}\leq(1-\psi_{jt\omega}^{C_{min}})M^{c},\quad\forall j,\forall t,\forall\omega$ (3.7g)
$\Delta_{j}^{C_{max}}-\Delta_{jt\omega}^{C}\leq(1-\psi_{jt\omega}^{C_{max}})M^{c},\quad\forall j,\forall t,\forall\omega$ (3.7h)
$q_{jt\omega}+\Delta_{jt\omega}^{C}\leq(1-\psi_{jt\omega})M^{c}\quad\forall j,\forall t,\forall\omega$ (3.7i)
$\nu_{jt\omega}^{C_{min}}\leq\psi_{jt\omega}^{C_{min}}M^{P},\quad\forall j,\forall t,\forall\omega$ (3.7j)
$\nu_{jt\omega}^{C_{max}}\leq\psi_{jt\omega}^{C_{max}}M^{P},\quad\forall j,\forall t,\forall\omega$ (3.7k)
$\varepsilon_{jt\omega}\leq\psi_{jt\omega}M^{P},\quad\forall j,\forall t,\forall\omega$ (3.7l)
$\psi_{jt\omega},\psi_{jt\omega}^{C_{min}},\psi_{jt\omega}^{C_{max}}\in\{0,1\}\quad\forall j,\forall t,\forall\omega$ (3.7m)
where $M^{c}$ and $M^{P}$ are large enough constants.
((3.5f)) and ((3.5g)) are linearized as follows:
$y_{t\omega}-\delta_{t\omega}\geq 0,\quad\forall t,\forall\omega$ (3.8a)
$y_{t\omega}+\delta_{t\omega}\geq 0,\quad\forall t,\forall\omega$ (3.8b)
$\alpha_{t\omega}^{+}\geq 0,\quad\forall t,\forall\omega$ (3.8c)
$\alpha_{t\omega}^{-}\geq 0,\quad\forall t,\forall\omega$ (3.8d)
$y_{t\omega}-\delta_{t\omega}\leq(1-\tau_{t\omega}^{+})M^{S}\quad\forall t,\forall\omega$ (3.8e)
$y_{t\omega}+\delta_{t\omega}\leq(1-\tau_{t\omega}^{-})M^{S}\quad\forall t,\forall\omega$ (3.8f)
$\alpha_{t\omega}^{+}\leq\tau_{t\omega}^{+}M^{T},\quad\forall t,\forall\omega$ (3.8g)
$\alpha_{t\omega}^{-}\leq\tau_{t\omega}^{-}M^{T},\quad\forall t,\forall\omega$ (3.8h)
$\tau_{t\omega}^{+},\tau_{t\omega}^{-}\in\{0,1\}$ (3.8i)
where $M^{S}$ and $M^{T}$ are large enough constants.
Finally, the linearization of ((3.5h))-((3.5k)) renders:
$q_{t\omega}^{S}\geq 0\quad\forall t,\forall\omega$ (3.9a)
$y_{t\omega}\geq 0\quad\forall t,\forall\omega$ (3.9b)
$\theta_{t\omega}\geq 0\quad\forall t,\forall\omega$ (3.9c)
$\beta_{t\omega}\geq 0\quad\forall t,\forall\omega$ (3.9d)
$q_{t\omega}^{S}\leq(1-\tau_{t\omega}^{q})M^{S},\quad\forall t,\forall\omega$ (3.9e)
$y_{t\omega}\leq(1-\tau_{t\omega}^{y})M^{S},\quad\forall t,\forall\omega$ (3.9f)
$\theta_{t\omega}\leq\tau_{t\omega}^{q}M^{T},\quad\forall t,\forall\omega$ (3.9g)
$\beta_{t\omega}\leq\tau_{t\omega}^{y}M^{T},\quad\forall t,\forall\omega$ (3.9h)
$\tau_{t\omega}^{q},\tau_{t\omega}^{y},\tau_{t}^{P}\in\{0,1\}$ (3.9i)
where $M^{S}$ and $M^{T}$ are large enough constants.
4 Numerical results and discussion
In this section, we provide numerical examples to complement the analytical results in the previous sections.
4.1 Data
We consider a day-ahead electricity market clearing price from the European Energy Exchange (EEX). The day-ahead electricity market clearing prices are available in one-hour intervals and can be found at EEX (2023), along with other market data. The marginal utility is the price-demand curve where $A_{jt\omega}$ and $B_{jt\omega}$ are price-demand curve intercept and slope, respectively. By utilizing day-ahead clearing prices and quantities data from a specific date, we can calibrate the expected values for utility parameters ($A_{jt\omega}$ and $B_{jt\omega}$) for consumer $j$, at time $t$ under scenario $\omega$. Subsequently, introducing perturbations allows for the consideration of varying levels of flexibility among residential consumers (Zugno et al., 2013). Regarding consumers’ flexibility, $\Delta_{j}^{C_{max}}=0$ indicates no flexibility, whereas $\Delta_{j}^{C_{max}}>0$ signifies demand flexibility of varying degrees.
Although theoretically there is no limit to the number of consumers, we simplify our model by limiting it to three. This limitation to three consumers is sufficient to yield quantitative insights into the model’s behavior because i) the findings can be generalized to representative real-world scenarios, as previously exemplified by Niromandfam et al. (2020) and Zugno et al. (2013). ii) we address the preferences of three consumers by adjusting demand parameters and offering flexibility in response to changes in consumption. Technically, this entails analyzing the scenarios for each consumer both before and after the implementation of the demand response program, effectively treating the three consumers as if we were dealing with six separate consumers. That means, the expected values of parameters $A_{j}$, $B_{j}$, and $\Delta_{j}^{C_{max}}$ are assumed to vary. This variation reflects the diverse reactions of consumers to price signals from the retailer. Furthermore, each consumer reflects distinct behavior such as Consumer 1 demonstrates the lowest slope and greatest flexibility, while Consumer 3 shows the least flexibility, and Consumer 2 lies between these two extremes (as detailed in Table 2). This approach not only aids in understanding the impact of these parameters on market outcomes but also enhances the visualization of the impact of each parameter on each consumer’s market outcomes.
[…]
Parameters
Cases $A_{1}$ $A_{2}$ $A_{3}$ $B_{1}$ $B_{2}$ $B_{3}$ $\Delta_{1}^{C_{max}}$ $\Delta_{2}^{C_{max}}$ $\Delta_{3}^{C_{max}}$
Benchmark 0.0291 0.0302 0.0271 0.0013 0.0015 0.0014 2.50 1.40 2.00
Linear 0.035 0.0375 0.0341 0.0013 0.0015 0.0014 2.50 1.40 2.00
Quadratic 0.0291 0.0302 0.0271 0.0017 0.020 0.0019 2.50 1.40 2.00
Flexibility 0.0291 0.0302 0.0271 0.0013 0.0015 0.0014 5.00 2.40 3.50
No Flexibility 0.0291 0.0302 0.0271 0.0017 0.020 0.0019 0.00 0.00 0.00
Linear utility 0.0291 0.0302 0.0271 0.00 0.00 0.00 5.00 2.40 3.50
4.2 Case studies
We consider six cases for both the MPEC and MILP (equilibrium) models in the numerical simulation:
  • i)
    Benchmark: This case serves as a reference for the increase or decrease of demand parameters ($A_{jt\omega},B_{jt\omega},\Delta_{j}^{C_{max}}$) calibrated using the input data from Table 2.
  • ii)
    Linear: Parameter $A_{j}$ is adjusted by 25% increase compared to its value in the Benchmark case, while keeping other parameters constant.
  • iii)
    Quadratic: Parameter $B_{j}$ is adjusted by 35% increase compared to its value in the Benchmark while keeping other parameters constant.
  • iv)
    Flexibility: Parameter $\Delta_{j}^{C_{max}}$ is adjusted by 82% increase compared to its value in the Benchmark case while keeping other parameters constant.
  • v)
    No flexibility: Parameter $\Delta_{j}^{C_{max}}$ is set to 00.
  • vi)
    Linear utility: parameter $B_{jt\omega}$ is set to 00 and the other parameters kept constant to the Benchmark values.
Based on these six cases, we illustrate the market outcomes for the proposed models, comparing the results across these case studies.
To accommodate the stochastic nature of input parameters in our model, we implement stochastic programming. This approach involves considering scenarios that are assumed to be equiprobable and applying a Gaussian distribution to each parameter. We then utilize the mean values in Table 2 to calibrate simulation data, ensuring the accuracy and reliability of our results. To this end, we select coefficients of variation (CVs) to determine the corresponding standard deviations. For instance, to compute the standard deviation of spot prices, we use the formula $CV\times E[P_{t\omega}^{S}]$, where $E[P_{t\omega}^{S}]$ represents the expected spot price at time $t$ under scenario $\omega$, and we have set the CV at $0.015$ for spot market calculation. Here, $E[P_{t\omega}^{S}]$ represents the day-ahead clearing prices for 24 hours under scenario $\omega$. Following this, $P_{t\omega}^{S}$—the spot price under scenario $\omega$—is used in the simulation to reflect $\omega$ possible scenario regarding market clearing prices. Similarly, we generate $A_{jt\omega}\sim N(\mu_{A_{jt\omega}},\sigma_{A_{jt\omega}})$ and $B_{jt\omega}\sim N(\mu_{B_{jt\omega}},\sigma_{B_{jt\omega}})$, where their standard deviations are calculated using CV values of 0.013 and 0.0013, respectively. Finally, we allow the demand flexibility parameter $\Delta_{j}^{C_{max}}$ to vary from 0 to 5.0 kWh. This range is chosen as it reflects realistic levels of flexibility for individual consumers participating in DR programs, as supported by our analysis and existing literature.
The models have been solved using JuMP version 1.11.1 (Lubin et al., 2023) under the open-source Julia programming Language version 1.8.5 (Bezanson et al., 2017). We use Artelys KNITRO solver version 13.2 (Artelys KNITRO, 2021) for the MPEC and Gurobi version 1.0.1 (Gurobi Optimization, 2021) for the MILP on a CPU E5-1650v2@3.50GHz and 64.00 GB of RAM running workstation. For the MPEC problem, we test the case study with $\Omega=30$ scenarios, and for the MILP problem $\Omega=300$. However, in the MPEC model, the number of scenarios increases, and the number of variables and constraints in the MPEC problem grows exponentially. This growth can potentially render the problem computationally intractable, especially when the number of scenarios exceeds 30($\Omega>30$) [^2] We do not present computational considerations by varying the number of scenarios for the sake of space.. There are possible alternatives to tackle this computational challenge such as utilizing high-performance computing (HPC) with parallel computing techniques (we refer the interested reader to Ahmadi-Khatir et al. (2013) and Nasri et al. (2015)). In our case, we employed the option: KNITRO multistart in parallel with 12 threads. This strategy discards low-quality solutions and initiates the algorithm at different random locations to select the best solutions, should any other local solutions be found. Moreover, a sensitivity analysis on the number of scenarios shows that increasing the number of scenarios has no significant impact on the reported market outcomes (see Table 3). It should also be noted that in the case of the MILP, the computational issue is considerably less concerning. For instance, the model can be efficiently solved even for larger sets of scenarios, such as $\Omega=500$ or greater. The MPEC takes $42.422$ CPU time with $\omega=30$, while MILP takes $15.019$ CPU time with $\omega=300$. It is also important to note that the methods to adjust big Ms are mostly heuristics and problem-dependent, where the trial-and-error procedure has been used in most research work (Motto et al., 2005; Garcés et al., 2009; Jenabi et al., 2013). However, for the particular problem addressed, the tune-up of these constants is achieved without particular numerical trouble. Nevertheless, caution must be exercised for the big M values related to the dual variables to ensure that their value is sufficiently large to not impose any additional constraints to the problem, but not too large so that the resulting model is not ill-conditioned (Pineda and Morales, 2019).
Table 3: Market outcomes sensitivity on the number of scenarios in the MPEC model under the Benchmark case.
Market outcomes Market outcomes
$\Omega$ Expected profit $P_{t}$ $q_{t\omega}^{S}$ $q_{jt\omega}$ $\Omega$ Expected profit $P_{t}$ $q_{t\omega}^{S}$ $q_{jt\omega}$
10 0.1857 0.0268 14.047 4.682 20 0.1839 0.0267 14.211 4.737
15 0.1850 0.0268 14.074 4.691 30 0.1844 0.02267 14.201 4.733
4.3 Market outcomes: Electricity price tariffs
The retailer’s price tariffs in both the MPEC and the MILP models are depicted along with the expected day-ahead electricity market prices in Figure 3b(a). The price tariffs are compared across the Benchmark, Flexibility, and No flexibility cases in both models. In all cases, the MILP price tariffs demonstrate consistent patterns with the market clearing prices in the day-ahead market. However, with the MPEC model, when the retailer’s market power is considered, price tariffs exceed the expected day-ahead electricity prices even during low-demand periods. This is market inefficiency, primarily due to market distortion caused by the strategy of the retailer.
(a) Comparison of price tariffs in both models across various cases and the day-ahead clearing prices.
(b) Variation in price tariffs with MPEC by adjusting model parameters.
Figure 3: Comparison of electricity price tariffs: MPEC vs. MILP across various cases.
On the other hand, within the competitive equilibrium model (MILP), price tariffs are generally observed to be lower than those in the MPEC model across most cases and periods. The overall price tariffs $P_{t}$ set by the retailer under the MPEC framework are consistently higher than the spot market prices $P_{t\omega}^{S}$. Furthermore, the retailer reduces its procurement when the spot market prices are relatively high (at $9:00,10:00,17:00-21:00$). On the other hand, it increases its purchases when the spot market prices are lower (at $t=1:00-8:00,10:00-17:00$). These electricity price patterns align with the quantities of power the retailer purchases during peak periods, which is reflected in the optimal outcomes showing lower quantities in the spot market. A more detailed analysis can be found in Figure 4, where there is a comprehensive comparison of the results both before and after the implementation of the demand response program under MPEC and MILP models. Higher quantities are purchased when wholesale market prices are lower. Although lower electricity price tariffs are already in place during the early hours, the introduction of increased demand flexibility results in a further decrease in electricity price tariffs in the MPEC model. However, in the equilibrium model (MILP), the overall electricity price tariff consistently remains below the spot price, except during peak periods.
The rise in consumer flexibility leads to lower retailer price tariffs in the MPEC, especially after 7:00. This indicates that if consumers adjust their consumption, electricity prices can decrease, and conversely, prices can increase with less flexibility, even in the presence of market distortion caused by strategic retailers. More interestingly, the MPEC model also has the peak-shaving/valley-filling effect; that is, the realized price tariffs at on-peak and off-peak periods are very similar in the case of high flexibility. This is a much-desired result, as a flatter price tariffs curve means that the system is more predictable and reliable. In the MILP model, however, high demand elasticity notably reduces electricity prices and increases the quantities purchased in both spot and retail markets to a greater extent than observed in the MPEC model. Thus, without demand flexibility, price tariffs are higher in the Benchmark under the MPEC model than in the MILP. Prices do not significantly change over time in the Benchmark case compared with case Flexibility case.
In Figure 3b(b), the impact of increasing demand parameters on price tariffs is illustrated. While an increase in $A_{jt\omega}$ slightly increases the clearing price, the overall impact of demand parameters and the penalty weight on price tariffs is insignificant. Note that imposing a penalty on the retailer’s objective function does affect the power imbalance in the periods considered rather than its impact on electricity price tariffs. Interestingly, increasing the flexibility parameter $\Delta_{j}^{C_{max}}$ leads to a significant reduction in price tariffs, more so than other model parameters. Upon analyzing the market outcomes of the proposed models, a key insight is that market power hinders the efficiency of demand flexibility, where the price-maker retailer exploits consumers’ flexibility to maximize its expected profits by imposing higher tariffs.
4.4 Market outcomes: Loads and spot market quantities
(a) Comparison of purchased quantities in both models before and after the implementation of the demand response program.
(b) Quantities purchased in spot market for both models before and after the implementation of the demand response program.
Figure 4: Comparison of consumption and spot market quantities with No flexibility and Flexibility cases.
Figure 4(a) depicts the aggregate consumer purchases in the retail market both before and after the implementation of demand response, in the MPEC and MILP models. The results indicate that lower spot market prices lead to increased purchases by consumers in the MILP model. The trend is similar without the implementation of the demand response program. With MILP, a higher value of $B$ (Quadratic case) is associated with significant and consistent electricity purchases in the retail market. Note that an increase in the quadratic term of the utility function is analogous to a decrease in electricity prices. Both parameters are inversely related to the quantity of electricity purchased. In contrast, an increase in the parameter in $A$ (Linear case) encourages consumers to buy more electricity during off-peak periods in both MPEC and MILP models, thereby avoiding higher tariff periods. Consequently, lower prices during off-peak periods prompt higher purchases from the spot and the retail markets. As loads shift from expensive to cheaper periods using consumers’ flexibility, new peak and valley periods emerge. This indicates that a peak period is typically followed by a valley period. Note that the detailed results for Linear and Quadratic cases are omitted for brevity.
With different cases, the retailer buys higher quantities in the perfect competition where there is a lower price, particularly with case Quadratic (higher $B$). With the Benchmark case, where there are lower values for $A$ and $\Delta_{j}^{C_{max}}$, the retailer purchases a lower quantity even in the MILP model. The equilibrium model shows the higher the competition, the lower the impact the retailer could have on the market outcomes. Specifically, the price tariffs sent by the retailer are lower in the MILP model as the retailer is not a price-maker. This enables consumers to enjoy greater benefits when purchasing their electricity in a competitive market model.
Figure 4(b) shows the quantities purchased in the spot market with MPEC and MILP under the Benchmark case. The spot market quantities show consistency in both models before and after implementing the DR program. However, with high Flexibility, the quantities in the spot market during off-peak periods are significantly higher. This shows that consumer flexibility allows retailers to shift their purchase plans to off-peak periods to benefit from lower prices in the wholesale market. During peak demand periods, on the other hand, retailers might opt out of buying from the day-ahead market due to the high spot market prices. If retailers have to purchase electricity during peak periods when wholesale prices are high, they will impose higher price tariffs on consumers during these periods. This, in turn, motivates consumers to shift/shave their consumption, which benefits both parties.
Despite the aggregate purchase trend reflecting individual consumer purchases, a comparison of consumers’ purchases in the Benchmark case reveals distinct consumer behaviors based on their utility function parameters. For instance, Consumer $1$ tends to buy more than other consumers when electricity price tariffs are low, typically in the early morning and late evening. This is because Consumer $1$ has the minimum slope ($A_{1tw}=0.0013$) and the highest demand flexibility ($\Delta_{j}^{C_{max}}=2.50$) in the Benchmark case. Individual consumers’ results are omitted for brevity.
(a) Price-quantity comparison: Benchmark.
(b) Price-quantity comparison: Linear.
(c) Price-quantity comparison: Quadratic.
(d) Price-quantity comparison: Flexibility.
Figure 5: Price (left axis) and spot quantity (right axis) comparison in both models and for various cases (Benchmark, Linear, Quadratic and Flexibiltiy).
4.5 Market outcomes: Electricity prices and retailer’s quantities comparison
Let’s expand the comparisons to include price tariffs, quantities, consumers’ utility parameters, and demand flexibility. Figure 5 presents the price tariffs sent by the retailer and quantities purchased by the retailer from the spot market in both models. Price tariffs are plotted on the left axis, and spot market quantities on the right axis. This allows for an analysis of how changes in utility parameters impact spot market purchases and price tariffs. Figure 5(a) depicts the Benchmark case, serving as a baseline for comparing the impacts of other parameters. In the quadratic utility function, the linear term (as seen in Figure 5 (b)) captures the initial positive impact of electricity consumption on consumers’ utility. Meanwhile, the Quadratic case models the diminishing returns in satisfaction as electricity consumption increases. As a result, when parameter $B_{jt\omega}$ increases, the retailer purchases a larger quantity from the wholesale market, as illustrated in Figure 5(c), and subsequently sells it back to consumers at lower price tariffs in the retail market. The lower price, in turn, increases consumer utility. It should be noted that from the retailer’s perspective, an increase in $B_{jt\omega}$ has an effect analogous to an increase in demand from the consumers.
Overall, the electricity price tariffs sent by the retailer and the quantities bought by the retailer from the spot market exhibit a similar trend. The proposed models effectively characterize the behaviors of market participants, facilitating the generalization of managerial and economic insights. Under the MILP model, consumers experience increased welfare due to lower prices compared to the MPEC model. Conversely, the MPEC model benefits the retailer, who acts as a price-maker, by enhancing its expected profits. Additionally, while prices remain relatively stable in the MPEC model, the quantities purchased from the spot market are significantly affected by the flexibility of consumers. Most notably, high consumer flexibility significantly impacts the quantity of purchases in the spot market. There is an uptick in purchases, primarily driven by lower prices during off-peak periods in the wholesale and retailer markets. The key takeaway is that as consumers become more willing to shift their consumption to periods of lower demand, the retailer has to buy from the wholesale market to fulfill demands during off-demand periods. This shift leads to an overall decrease in wholesale electricity prices and also enhances overall social welfare as detailed in Figure 7. From the perspective of market regulators, an important implication is that, depending on the market structure, tariffs-based incentives can have a significant impact on both spot and retail market outcomes.
4.5.1 Market outcomes: expected profit
Figure 6: Comparison of expected profit between two models before and after DR implementation.
The expected profit results of the retailer are depicted in Figure 6 in both models. The expected profit results are directly related to the price tariffs and quantities previously discussed. In the MILP model, the retailer’s myopic approach to market information results in financial losses in specific periods. This occurs because the model is specifically designed to maximize the expected total profit over an entire 24-hour time horizon [^3] The underlying assumption here is that consumers commit to a tariff agreement with the retailer, which is expected to remain valid for the specified time horizon. This approach is both reasonable and practical in the context of current electricity markets, as discussed in (Wu et al., 2015).. However, the consumers’ flexibility with DR plays a pivotal role, in enhancing the retailer’s expected profits and mitigating monetary losses. Notably, in the MPEC models, where the retailer adopts a strategic approach, its expected profits are significantly higher.
4.5.2 Market outcomes: Social welfare analysis
Social welfare (SW) is calculated as follows:
$SW$ $=\underbrace{\sum_{\omega\in\Omega}\sigma_{\omega}\left(\sum_{j\in J,t\in T}P_{t}q_{jt\omega}-\sum_{t\in T}(P_{t\omega}^{S}q_{t\omega}^{S}+Cy_{t\omega})\right)}_{\text{Retailer's expected profit}}+$
$\underbrace{\quad+\sum_{t\in T}\left(P_{t}q_{jt\omega}-A_{jt\omega}(q_{jt\omega}+\Delta_{jt\omega}^{C})+\frac{1}{2}B_{jt\omega}(q_{jt\omega}+\Delta_{jt\omega}^{C})^{2}\right)}_{\text{consumers' benefit minus cost of consumption}}.$
(a) Social welfare with a quadratic utility function.
(b) Social welfare with a linear utility function.
Figure 7: Social welfare for MPEC and equilibrium models in the No flexibility and Flexibility cases.
Figure 7 depicts social welfare with two models before and after the implementation of the DR program. Interestingly, the MPEC model explicitly highlights the social cost incurred when consumption is shifted to low-demand periods due to consumers’ flexibility. This suggests that price-maker retailers have the ability to manipulate this consumer flexibility, exploiting it to maximize their expected profits. In doing so, they transfer risks centered around the retailer to the consumers. On the other hand, it also demonstrates an increase in social welfare during peak-demand periods because consumers have the choice to adjust their consumption during these periods using dynamic price tariffs. Social welfare is better during low-demand periods due to low-price tariffs, and worse during peak-demand periods due to high-price tariffs before DR in both models. Despite the retailer’s exploitation, consumers’ flexibility is crucial in enhancing social welfare, especially during extended periods from 9:00 to 22:00. These hours are typically stressful for system operators, underlining the significance of consumer flexibility for system stability. In contrast, the equilibrium model exhibits positive social welfare in all periods. There are higher welfare gains during peak-demand periods, and the social cost during low-demand periods is significantly lower than that of the MPEC model. More importantly, with Figure 7(b), we present an analysis of social welfare before and after implementing the demand response program, using a Linear utility function. The findings from both models reveal that social welfare significantly improves with a linear utility function compared to a quadratic one. Interestingly, during off-peak periods, social welfare appears to be better without flexibility. However, despite the differences in the magnitude of welfare, the trend remains consistent whether using linear or quadratic utility functions. This consistency enables the generalization of market outcomes across the models, regardless of the specific functional form of the consumers’ utility functions.
(a) Consumers’ utility for each consumer: MPEC.
(b) Consumers’ welfare for each consumer: MPEC.
(c) Consumers’ utility for each consumer: MILP.
(d) Consumers’ welfare for each consumer: MILP.
Figure 8: Comparison of individual consumers’ utility and welfare in the MPEC and MILP using No flexibility and Flexibility cases.
Figure 8 shows individual consumers’ welfare in both models, before and after implementing the DR program. Consumers generally experience higher surpluses, especially during low-demand periods by using MILP. Consumers’ flexibility leads to a surplus for consumers across all periods compared to their utility without the DR program. Even with the MPEC model, consumers offset their losses during low-demand in the time periods from 8:00 to 22:00. Note that the differences in individual consumers’ surplus are attributable to variations in utility and demand flexibility parameters, as detailed in Table 2.
Figure 9 presents results concerning consumers’ flexibility in their consumption for MPEC and MILP: the Benchmark and Flexibility cases. The overall trend for flexible demand is similar in both models, although there are differences in magnitude. Notably, consumers demonstrate increased flexibility from 9:00-11:00 and 17:00-22:00, shifting their consumption to off-peak periods. Thus, indicating higher energy consumption during these off-peak times. The results adhere to the constraint $\sum_{t\in{T}}\Delta_{jt\omega}^{C}=0$ and respect the flexibility boundary values for each consumer as presented in Table 2. For example, consumer 1 exhibits higher flexibility in consumption, up to 5kWh, compared to other consumers. With high flexibility, the overall electricity price decreases, and the quantity purchased increases, leading to robust impacts on both the retailer’s expected profit and consumer surplus. Therefore, the DR models effectively characterize the market outcomes in both market configurations, highlighting the importance of utility functions in understanding consumer behavior and market dynamics.
(a) Level of flexibility for consumers: Benchmark.
(b) Level of flexibility for consumers: Flexibility.
Figure 9: Variation of demand in both models (left MPEC and right MILP) with respect to the Benchmark and Flexibility cases.
4.5.3 Market performance and sensitivity analysis
Table 4: Market performance of retailer and consumers in the simulations with the Benchmark case. All values, except for the price (€/kWh), represent average outcomes for the considered scenarios ($\Omega=30$) and are expressed in €.
Retailer Model Consumers Model
Performance index MPEC MILP performance index MPEC MILP
Expected profit 0.1149 -0.0076 Cost 0.2602 0.2894
Revenue 0.2602 0.2894 Price 0.0230 0.0177
Cost 0.1454 0.2970 Welfare 0.0177 0.0784
Table 4 presents the average performance of the two models for both the retailer and consumers under the Benchmark case. On average, the retailer incurs losses in the MILP model. The retailer has two sources of costs: i), cost incurred from purchasing in the spot market, which is based on the assumption of perfect information about future consumption; and ii) imbalance penalties that arise from imperfect information. It is essential to understand that the revenue generated by the retailer translates into costs for consumers, which significantly affects the overall dynamics of the market.
Table 5: Sensitivity analysis of spot price uncertainty in the MPEC model: Benchmark case.
Spot price uncertainty Exp. profit Max. profit Min. profit $q_{t\omega}$ $q_{jt\omega}$ Tariff Utility Welfare
0.015 0.1143 0.3555 -0.0128 12.29 4.198 0.0230 0.1046 0.0177
0.030 0.1156 0.3551 -0.0140 12.30 4.101 0.0230 0.1046 0.0177
0.035 0.1156 0.3547 -0.0141 12.31 4.105 0.0230 0.1047 0.0178
In Table 5, we present the market outcomes regarding the uncertainty of spot market prices $P_{t\omega}^{S}$ under the Benchmark case and the MPEC model. On the contrary, consumers enjoy better utility due to the low price tariffs. We consider the coefficient of variations to be 0.015, 0.003, and 0.035. The results indicate that uncertainties in spot market prices do not significantly influence market outcomes for either retailers or consumers. Instead, as demonstrated by our simulation results, the most crucial parameter affecting market outcomes is demand flexibility. This flexibility in consumer demand is key not only in mitigating the potential impact on market outcomes but also in influencing spot market price formation. This highlights the pivotal role of consumer flexibility in managing market dynamics and aiding the integration of renewable energy sources.
5 Conclusions
This paper introduces two game-theoretical models to analyze retailer-consumer interactions in electricity markets under demand response programs. The paper model consumer flexibility under dynamic pricing and employs a utility function to characterize consumer behavior, providing insights into market dynamics and reshaping their interpretation.
Two market structures are considered: a leader-follower model using a bi-level problem solved as MPEC and a perfect competition model solved as a MILP problem. The MPEC model facilitates a comparison of market power dynamics, while the MILP model helps in understanding competitive market behavior. Both models reveal their respective impacts on market outcomes and social welfare. This work advances existing research in demand response by introducing complex dynamics under dynamic price tariffs. It achieves this by simultaneously analyzing consumers’ utility functions, incorporating uncertainty, and considering various levels of market structures both in the spot and retail markets. These multifaceted interactions enable a more nuanced and comprehensive understanding of market dynamics, potentially reshaping how these dynamics are modeled and interpreted.
The findings demonstrate the effectiveness of dynamic pricing in demand response programs, optimizing strategic retailers’ expected profit and maximizing consumers’ welfare. The models’ adaptability to various consumer groups and their flexibility to retailers’ pricing signals are crucial for demand-side management under demand response programs. Additionally, the research presents compelling economic incentives for flexible consumers under price uncertainties, benefiting both retailers and consumers and suggesting a symbiotic relationship between DR programs and electricity markets.
The proposed methodology delivers crucial managerial and economic insights: i) it manages both demand and supply uncertainties through stochastic programming. ii) the analysis explores market configurations for each participant, focusing on enhancing flexibility in consumption and load management through dynamic tariffs. This approach aims to optimize the market structure for the overall improvement of social welfare. We have demonstrated that consumers’ flexibility is crucial in devising market strategies that are attuned to individual needs and adapt to varying market conditions. iii) the results reveal that utilizing consumers’ flexibility under tariff-based demand response renders significant cost savings and economic efficiency in electricity markets. The takeaway is that these benefits are crucial in promoting the sustainability and resilience of future electricity markets, underlying that demand response programs are key tools in energy transition under various market structures.
The proposed game-theoretical framework presents various opportunities for advanced research and application. First, it is possible to extend this framework to include the operational dynamics of generating companies and market operators. This enhancement would elucidate the reciprocal influences and impacts between these entities and demand response programs. Second, the models could be extended accounting for generation from renewable sources and grid constraints explicitly, coupled with the simulation of consumer behavior facilitated by smart appliances. This integration would enable the investigation of consumer-centric energy trading scenarios, evolving the role of consumers from passive end-users to proactive market contributors. Third, in light of the scant research on the environmental ramifications of demand response programs, integrating environmental metrics, particularly pollution metrics, into the demand response models is proposed. This could provide robust policy guidance for harmonizing technological advancements, market mechanisms, and innovative approaches within the electricity markets.
Acknowledgements
Arega Getaneh Abate has received funding from the European Union’s Horizon $2020$ research and innovation program under the Marie Skłodowska-Curie grant agreement No $899987$. Carlos Ruiz gratefully acknowledges the financial support from the Spanish government through projects PID2020-116694GB-I00 and from the Madrid Government (Comunidad de Madrid) under the Multiannual Agreement with UC3M in the line of “Fostering Young Doctors Research” (ZEROGASPAIN-CMUC3M) and in the context of the V PRICIT (Regional Programme of Research and Technological Innovation.